Two point charges, QA = +18 μC and QB= -5 μC, are separated by a distance r= 10 cm. What is the magnitude of the electric force. The constant k= 8.988 x 109 Nm2C


1
Expert's answer
2021-06-22T09:47:56-0400

From Coulomb's law, we know that the magnitude of the electric force will be equal to


"F=k\frac{|q_Aq_B|}{r^2}"


Then, we substitute (qA=18 "\mu"C, qB=-5 "\mu"C, r=10 cm and k = 8.988 X 109 Nm2/C2) and we can calculate the magnitude of such force:


"F=(8.988\times10^9\,Nm^2/C^2)\large{\frac{|(18\mu C)(-5\mu C)|}{(10\,cm)^2}}"


"\implies F=\frac{(8.988\times10^9\,Nm^2/C^2)|(18\times10^{-6} C)(-5\times10^{-6} C)|}{(0.1\,m)^2}"

"\implies F=\frac{(8.988)(18)(5)}{\large{(10^{-1})^2}} \times10^{9-12}\,N"


"\implies F=80892 \times10^{-3}\,N=80.892 \,N"


In conclusion, the magnitude of the electric force between these two charges under the present conditions is approximately equal to 80.892 N.


Reference:

  • Serway, R. A., & Jewett, J. W. (2018). Physics for scientists and engineers. Cengage learning.